# JoyAvatar: Real-time and Infinite Audio-Driven Avatar Generation with Autoregressive Diffusion

### v1

@@ Abstract
Existing DiT-based audio-driven avatar generation methods have achieved considerable progress, yet their broader application is constrained by limitations such as high computational overhead and the inability to synthesize long-duration videos. Autoregressive methods address this problem by applying block-wise autoregressive diffusion methods. However, these methods suffer from the problem of error accumulation and quality degradation. To address this, we propose JoyAvatar, an audio-driven autoregressive model capable of real-time inference and infinite-length video generation with the following contributions: (1) Progressive Step Bootstrapping (PSB), which allocates more denoising steps to initial frames to stabilize generation and reduce error accumulation; (2) Motion Condition Injection (MCI), enhancing temporal coherence by injecting noise-corrupted previous frames as motion condition; and (3) Unbounded RoPE via Cache-Resetting (URCR), enabling infinite-length generation through dynamic positional encoding. Our 1.3B-parameter causal model achieves 16 FPS on a single GPU and achieves competitive results in visual quality, temporal consistency, and lip synchronization.
Figure 1: JoyAvatar achieves real-time, streaming audio-driven avatar video generation at 16 FPS on a single GPU, while supporting infinite-length video synthesis with minimal error accumulation.
@@ 1 Introduction
To address these challenges, we present JoyAvatar, a 1.3B-parameter, audio-driven model employing a 4-step denoising process within a causal architecture. Our model enables real-time, streaming generation with infinite length, audio-synchronized talking videos. It is designed to maintain high visual clarity and fidelity, ensure consistent identity preservation, and achieve robust temporal coherence throughout extended sequences.
@@ 3 Methods
Figure 2: Overview of JoyAvatar. Sub-step Denosing Blocks implement Progressive Step Bootstrapping (PSB), Condition Frames embody Motion Condition Injection (MCI), while Sink Frames connected via preceding arrows demonstrate the mechanism of Unbounded RoPE via Cache-Resetting (URCR).
@@ 4.2 Comparison
We evaluate JoyAvatar and its bidirectional teacher model against state-of-the-art audio-driven video generation methods. Quantitative results across key metrics are presented in Table 1.
@@
As shown in Table 1, our bidirectional teacher model achieves superior performance across all evaluated metrics—including video quality, lip‑sync accuracy, and identity preservation—compared to existing open‑source methods. When distilled into a causal variant via Diffusion Model Distillation (DMD), the resulting JoyAvatar model shows a moderate performance decline, which is attributable to its reduced parameter scale and the constraints of its causal architecture. Nonetheless, JoyAvatar maintains competitive performance across all metrics relative to bidirectional models of comparable scale. Notably, it significantly outperforms the 14B‑parameter causal model LiveAvatar [21] in terms of Sync‑C and Sync‑D scores, while delivering comparable results on IDC, Q_score, and Q_face metrics. Importantly, the causal design yields a substantial improvement in inference speed: the DiT component reaches 16.2 FPS on a single GPU. With multi‑GPU parallelization and streaming optimizations, the full pipeline exceeds 30 FPS, enabling real‑time display and interactive applications.

### v2

# JoyAvatar-Flash: Real-time and Infinite Audio-Driven Avatar Generation with Autoregressive Diffusion

@@ Abstract
Existing DiT-based audio-driven avatar generation methods have achieved considerable progress, yet their broader application is constrained by limitations such as high computational overhead and the inability to synthesize long-duration videos. Autoregressive methods address this problem by applying block-wise autoregressive diffusion methods. However, these methods suffer from the problem of error accumulation and quality degradation. To address this, we propose JoyAvatar-Flash, an audio-driven autoregressive model capable of real-time inference and infinite-length video generation with the following contributions: (1) Progressive Step Bootstrapping (PSB), which allocates more denoising steps to initial frames to stabilize generation and reduce error accumulation; (2) Motion Condition Injection (MCI), enhancing temporal coherence by injecting noise-corrupted previous frames as motion condition; and (3) Unbounded RoPE via Cache-Resetting (URCR), enabling infinite-length generation through dynamic positional encoding. Our 1.3B-parameter causal model achieves 16 FPS on a single GPU and achieves competitive results in visual quality, temporal consistency, and lip synchronization.
Figure 1: JoyAvatar-Flash achieves real-time, streaming audio-driven avatar video generation at 16 FPS on a single GPU, while supporting infinite-length video synthesis with minimal error accumulation.
@@ 1 Introduction
To address these challenges, we present JoyAvatar-Flash, a 1.3B-parameter, audio-driven model employing a 4-step denoising process within a causal architecture. Our model enables real-time, streaming generation with infinite length, audio-synchronized talking videos. It is designed to maintain high visual clarity and fidelity, ensure consistent identity preservation, and achieve robust temporal coherence throughout extended sequences.
@@ 3 Methods
Figure 2: Overview of JoyAvatar-Flash. Sub-step Denosing Blocks implement Progressive Step Bootstrapping (PSB), Condition Frames embody Motion Condition Injection (MCI), while Sink Frames connected via preceding arrows demonstrate the mechanism of Unbounded RoPE via Cache-Resetting (URCR).
@@ 4.2 Comparison
We evaluate JoyAvatar-Flash and its bidirectional teacher model against state-of-the-art audio-driven video generation methods. Quantitative results across key metrics are presented in Table 1.
@@
As shown in Table 1, our bidirectional teacher model achieves superior performance across all evaluated metrics—including video quality, lip‑sync accuracy, and identity preservation—compared to existing open‑source methods. When distilled into a causal variant via Diffusion Model Distillation (DMD), the resulting JoyAvatar-Flash model shows a moderate performance decline, which is attributable to its reduced parameter scale and the constraints of its causal architecture. Nonetheless, JoyAvatar-Flash maintains competitive performance across all metrics relative to bidirectional models of comparable scale. Notably, it significantly outperforms the 14B‑parameter causal model LiveAvatar [11] in terms of Sync‑C and Sync‑D scores, while delivering comparable results on IDC, Q_score, and Q_face metrics. Importantly, the causal design yields a substantial improvement in inference speed: the DiT component reaches 16.2 FPS on a single GPU. With multi‑GPU parallelization and streaming optimizations, the full pipeline exceeds 30 FPS, enabling real‑time display and interactive applications.